# references.bib

### v4

@@ Abstract
This research comprehensively describes the basic theory of transversally Heisenberg elliptic operators, and investigates the index theory of Heisenberg elliptic and transversally Heisenberg elliptic operators from the perspective of $KK$-theory, applying Kasparov’s methodology. Moreover, the analysis methodically examines specific conditions, with a focus on the Fourier transform of the nilpotent group $C^{\ast}$-algebra. We demonstrate enhanced methods for analyzing the hypoellipticity of operators, presenting a robust framework for defining and understanding transversal Heisenberg ellipticity in a $KK$-theoretic context. This work provides a solid foundation for future research into the properties of differential operators in complicated manifolds.
@@ Introduction
Let $M$ be a smooth and closed manifold, $E,F\to M$ be vector bundles, and $P:C^{\infty}(M;E)\to C^{\infty}(M;F)$ be a differential operator over $M$. For elliptic operators, the spaces $\text{Ker}(P)$ and $\text{Ker}(P^{\ast})$ are both finite-dimensional, and the analytical index is defined as $\text{Ind}_{a}(P)=\dim\text{Ker}(P)-\dim\text{Coker}(P)\in\mathbb{Z}$. When a Lie group $G$ acts on $M$ and the operator $P$ is elliptic, the spaces $\text{Ker}(P)$ and $\text{Coker}(P)$ can be viewed as elements of $R(G)$, with the analytical index defined as $\text{Ind}_{a}(P)=[\text{Ker}(P)]-[\text{Coker}(P)]\in R(G)$. The Atiyah-Singer index theorem [atiyah1968index2][atiyah1968index1] states that in each case, the analytical index defined above is equal to the topological index, which only depends on the topological characteristics of $P$ and $M$.
@@
Kasparov [kasparov2017elliptic][kasparov2022k] generalized it for locally compact manifolds under properly supported Lie group actions, framing a comprehensive theory on the index of elliptic and transversally elliptic operators through $KK$-theory. In this theory both the symbol class $[\sigma(P)]$ and the index class $[P]$ are represented by $KK$-classes, providing the index theorem through $KK$-products.
@@
On a closed manifold with filtration, or equivalently on a closed manifold endowed with a regular foliation, the index theorem for Heisenberg elliptic (later on, H-elliptic) operators can be reduced as a special case of Mohsen’s maximally hypoelliptic index theorem. In the regular foliation case, the characteristic set defined in [androulidakis2022pseudodifferential] is the entire groupoid $Gr(H)$. Hence, the $\ast$-maximally hypoelliptic condition is consistent with the H-ellipticity, which requires for each nontrivial unitary irreducible representation $\pi$, the action on the H-cosymbol $\sigma_{H}(P,x,\pi)$ defined by the filtration is bijective. We will also provide an index theorem for H-elliptic operators from the $KK$-theoretic viewpoint in Section 3.2.
This research introduces transversally H-elliptic operators on filtered manifolds under a $G$-action, a concept that is inspired by Kasparov’s tangent-Clifford symbol class $[\sigma^{tcl}(P)]$ in transversally elliptic index theory [kasparov2022k]. Transversal H-ellipticity requires the Rockland condition’s fulfillment in conjunction with a leafwise elliptic symbol class. Concretely, We define that:
@@
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group acting on $M$ isometrically and preserving the filtered structure. A pseudo-differential operator $P$ of 00-order is called transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in M(E_{x}\otimes_{\mathbb{C}}C^{\ast}(Gr(H)_{x})),$
@@
In Chapter 4, We will prove that Definition 0.0.1 produces an appropriate tangent-Clifford symbol class $[\sigma^{tcl}_{H}(P)]$ and an index class $[P]$ for each transversally H-elliptic operator $P$. When $P_{0}$ is a transversally H-elliptic differential operator, let $P$ be the bounded transform of $P_{0}$. Then, there is an index formula in $KK$-homology in which $[\sigma^{tcl}_{H}(P)]$ and $[P]$ can be related using $KK$-products. Here is the main theorem of the paper:
@@
Let $M$ be a closed filtered Riemannian manifold and $G$ be a Lie group isometrically acting on $M$ and preserving the filtered structure. Let $P_{0}$ be a $G$-invariant differential operator on $M$ with its bounded transform
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix}$
being transversally H-elliptic of order 00. Then, we have the following formula
$[P]=j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H)))}j^{G}([\mathcal{E}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}],$
where $[\sigma^{tcl}_{H}(P)]\in K_{0}^{G}(C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$ is the tangent-Clifford H-cosymbol for $P$, $[\mathcal{E}]\in KK_{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$ refers to the Connes-Thom class and $[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}]\in K^{0}(C^{\ast}(G,$ $Cl_{\tilde{\Gamma}}(TM)))$ indicates the Clifford–Dolbeault element.
Determining whether an arbitrary pseudo-differential operator meets the criteria outlined in our defined theory hinges on the Fourier transform theory for nilpotent group $C^{\ast}$-algebras. For an element $\phi\in C^{\ast}(\mathcal{G})$, where $\mathcal{G}$ represents a nilpotent Lie group, its representations $\pi(\phi)$ constitute an operator-valued function on the spectrum $\widehat{\mathcal{G}}$, with some specific conditions. This theory extends to the Fourier transform of multiplier elements of the group $C^{\ast}$-algebra, transforming a multiplier of $C^{\ast}(\mathcal{G})$ into an operator family parametrized by $\widehat{\mathcal{G}}$. As a result, a multiplier of $C^{\ast}(\mathcal{G})$ is Fourier-transformed to a family of operators parametrized by the representation space $\widehat{\mathcal{G}}$, and the properties of this operator family reflect if the multiplier is an element of $C^{\ast}(\mathcal{G})$ itself or not. Explicitly, for an element $e\in M(C^{\ast}(\mathcal{G}))$, if we can verify that the Fourier transform $\mathcal{F}(e)\in\mathcal{F}(M(C^{\ast}(\mathcal{G})))$ actually lies in $\mathcal{F}(C^{\ast}(\mathcal{G}))$, then $e\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$. See the following diagram:
@@
For each $x\in M$, suppose that $Q$ is a multiplier of the $C^{\ast}(Gr(H)_{x})$ module, $Q\in M(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$. Then, $Q$ is compact , i.e.,
@@
lies in the sub-algebra $\text{End}(E_{x})\otimes B^{\ast}(\widehat{Gr(H)}_{x})$. The algebra $B^{\ast}(S)$ is defined in Definition 6.1.5.
An illustrative example of transversally H-elliptic operators is given by the operator $P_{\gamma}=\Delta_{n-k}+i\gamma T$ on the Heisenberg group $M=\mathbb{H}_{n}$ with transition action by $\mathbb{R}^{2k}$ on the first $2k$-coordinates. Here, $\Delta_{n-k}=-\sum_{i=k+1}^{n}({X_{i}}^{2}+Y_{i}^{2})$ means the sub-Laplacian operator on the sub-manifold $M^{\prime}\cong\mathbb{H}_{n-k}$ identified by coordinates $x_{k+1},y_{k+1},\dots,x_{n},y_{n},z$. The function $\gamma$ on $M$ is assigned such that $P_{\gamma}$ on $M^{\prime}$ is H-elliptic, see [baum2014k]. Utilizing Theorem 0.0.3, we can show that $P_{\gamma}$ is transversal H-elliptic on $M$.
@@
  • Chapter 3 articulates the general index theorem for H-elliptic operators, extending van Erp’s findings on contact manifolds [van2010atiyah1] and presenting the theorem from a $KK$-theoretic perspective.
@@
  • Chapter 5 is dedicated to the main index theorem for transversally H-elliptic operators. We give a proof for the H-elliptic index theorem in Chapter 3 and extend it to transversally H-elliptic operators.
  • Chapter 6 introduces a Fourier transform theory for $C^{\ast}$-algebra of nilpotent Lie groupoids [beltictua2017fourier] to verify the compactness condition of Definition 0.0.1.
  • In the Appendix, we exhibit other possible conditions to define the transversally H-elliptic operators. We include prototypes of these definitions, from the definitions or certain crucial properties of transversally elliptic or H-elliptic operators.
@@
  • We denote by $G$ the Lie group act on manifold $M$, by $\mathcal{G}$ a groupoid or a bundle of nilpotent Lie groups over the manifold $M$.
@@
  • If $E$ is a vector bundle over $M$, we will denote by $C^{\infty}(M;E)$ the space of smooth sections of $E$, $C(M;E)$ the space of continuous sections of $E$ and $C_{0}(M;E)$ the space of continuous sections of $E$ vanishing at infinity. We denote $C_{0}(E)$ to mean the space of continuous functions on $E$ (a topological space) vanishing at infinity.
@@
  • We denote by $M(A)$ the multiplier algebra for a $C^{\ast}$-algebra $A$.
@@
In this chapter, we introduce fundamental concepts related to filtered manifolds and the symbolic calculus of operators on these manifolds.
@@ 1.1 Filtered manifold and the osculating groupoid
Let $M$ be a smooth closed manifold with a filtration of the tangent bundle
@@
Assume that $M$ is equipped with a proper isometric action by a locally compact group $G$.
@@
At each point $x$ on the filtered manifold $M$, there is a simply-connected nilpotent Lie group structure on $T_{x}M$, called the osculating group at $x$.
@@
It can be proved that $[X,Y](x)$ mod $H^{i+j-1}_{x}$ depends only on $X(x)$ mod $H^{i-1}_{x}$ and $Y(x)$ mod $H^{j-1}_{x}$. Therefore, the Lie bracket structure over $gr(H)_{x}$ is well defined for each $x\in M$. The graded algebra $gr(H)_{x}$ at each point is called the osculating Lie algebra. Then there is a corresponding simply-connected nilpotent Lie group for each $x$, we denote it by $Gr(H)_{x}$. ∎
@@ 1.2 Kirillov’s orbit method
The primary goal of Kirillov’s orbit method, as outlined in [kirillov2004lectures], is to characterize the unitary dual $\widehat{G}$ as a topological space and to construct the unitary irreducible representations $\pi_{\Omega}$ associated with a coadjoint orbit $\Omega\in\mathfrak{g}^{\ast}$.
@@
Kirillov’s orbit method asserts the following.
@@
This map from coadjoint orbits $\Omega\in\mathcal{O}(G)$ in $\mathfrak{g}^{\ast}$ to unitary irreducible representations of $\pi_{\Omega}\in\widehat{G}$ establishes a one-to-one correspondence.
@@ 1.3 Groupoid $C^{\ast}$-algebra
      sourse and range maps $r,s:\mathcal{G}\rightrightarrows\mathcal{G}^{(0)}$;
@@
      a multiplication map $\mathcal{G}^{(2)}\to\mathcal{G}$, $(a,b)\to ab$, where
@@
      an inverse map $\mathcal{G}\to\mathcal{G}$, $a\to a^{-1}$, such that:

### v5

@@ Abstract
This research comprehensively describes the basic theory of transversally Heisenberg elliptic operators, and investigates the index theory of Heisenberg elliptic and transversally Heisenberg elliptic operators from the perspective of $KK$-theory, applying Kasparov’s methodology. Moreover, the analysis methodically examines specific conditions, with a focus on the Fourier transform of the nilpotent group $C^{\ast}$-algebra. We demonstrate enhanced methods for analyzing the hypoellipticity of operators, presenting a robust framework for defining and understanding transversal Heisenberg ellipticity in a $KK$-theoretic context. This work provides a solid foundation for future research into the properties of hypoelliptic differential operators in complicated manifolds.
@@ Introduction
Let $M$ be a smooth and closed manifold, $E,F\to M$ be vector bundles, and $P:C^{\infty}(M;E)\to C^{\infty}(M;F)$ be a differential operator over $M$. For elliptic operators, the spaces $\text{Ker}(P)$ and $\text{Ker}(P^{\ast})$ are both finite-dimensional, and the analytical index is defined as $\text{Ind}_{a}(P)=\dim\text{Ker}(P)-\dim\text{Coker}(P)\in\mathbb{Z}$. When a compact Lie group $G$ acts on $M$ and the operator $P$ is elliptic, the spaces $\text{Ker}(P)$ and $\text{Coker}(P)$ can be viewed as elements of $R(G)$, with the analytical index defined as $\text{Ind}_{a}(P)=[\text{Ker}(P)]-[\text{Coker}(P)]\in R(G)$. The Atiyah-Singer index theorem [atiyah1968index2][atiyah1968index1] states that in each case, the analytical index defined above is equal to the topological index, which only depends on the topological characteristics of $P$ and $M$.
@@
Kasparov [kasparov2017elliptic][kasparov2022k] generalized it for locally compact manifolds under properly supported locally compact Lie group actions, framing a comprehensive theory on the index of elliptic and transversally elliptic operators through $KK$-theory. In this theory both the symbol class $[\sigma(P)]$ and the index class $[P]$ are represented by $KK$-classes, providing the index theorem through $KK$-products.
@@
On a closed manifold with filtration, or equivalently on a closed manifold endowed with a regular foliation, the index theorem for Heisenberg elliptic (later on, H-elliptic) operators can be reduced as a special case of Mohsen’s maximally hypoelliptic index theorem. In the regular foliation case, the characteristic set defined in [androulidakis2022pseudodifferential] is the entire groupoid $Gr(H)$. Hence, the $\ast$-maximally hypoelliptic condition is consistent with the H-ellipticity, which requires for each nontrivial unitary irreducible representation $\pi$, the action on the H-cosymbol $\sigma_{H}(P,x,\pi)$ defined by the filtration is bijective. We will reformulate the index theorem for H-elliptic operators from the $KK$-theoretic viewpoint in Section 3.2 and 3.3.
In this research, we introduce transversally H-elliptic operators on filtered manifolds with $G$-action. The definition is inspired by Kasparov’s tangent-Clifford symbol class $[\sigma^{tcl}(P)]$ in transversally elliptic index theory [kasparov2022k]. Transversal H-ellipticity requires the Rockland condition’s fulfillment in conjunction with a leafwise elliptic symbol class. Concretely, We define that:
@@
Let $M$ be a complete filtered Riemannian manifold and $G$ be a locally compact Lie group acting on $M$ properly, isometrically, and preserving the filtered structure. A pseudo-differential operator $P$ of 00-order is called transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in\mathcal{M}(E_{x}\otimes C^{\ast}(Gr(H)_{x})),$
@@
In Chapter 4, We will prove that Definition 0.0.1 produces an appropriate tangent-Clifford symbol class $[\sigma^{tcl}_{H}(P)]$ and an index class $[P]$ for each transversally H-elliptic operator $P$. When $P$ is a transversally H-elliptic operator of 00-order, there is an index formula in $KK$-homology in which $[\sigma^{tcl}_{H}(P)]$ and $[P]$ can be related using $KK$-products. Here is the main theorem of this paper:
@@
Let $M$ be a complete filtered Riemannian manifold and $G$ be a locally compact Lie group isometrically and properly acting on $M$ and preserving the filtered structure. Let $P$ be a $G$-invariant transversally H-elliptic operator on $M$ of order 00, then we have
$[P]=j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M))}[\mathcal{D}_{M,{\Gamma}}^{cl^{\prime}}],$
in which $[\mathcal{D}_{M,{\Gamma}}^{cl^{\prime}}]\in K^{0}(C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M)))$ indicates the Clifford-Dolbeault element which will be defined in Chapter 5, and the $KK$-product on the right-hand side is considered over
$KK(C^{\ast}(G,C_{0}(M)),C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{{\Gamma}}(M))$
$\otimes_{C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M))}KK(C^{\ast}(G,C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{\Gamma}(M)),\mathbb{C}).$
Determining whether an arbitrary pseudo-differential operator meets the criteria outlined in our defined theory hinges on the Fourier transform theory for nilpotent group $C^{\ast}$-algebras. For an element $\phi\in C^{\ast}(\mathcal{G})$, where $\mathcal{G}$ represents a nilpotent Lie group, its representations $\pi(\phi)$ constitute an operator-valued function on the spectrum $\widehat{\mathcal{G}}$, with some specific conditions. This theory extends to the Fourier transform of multiplier elements of the group $C^{\ast}$-algebra, transforming a multiplier of $C^{\ast}(\mathcal{G})$ into an operator family parametrized by $\widehat{\mathcal{G}}$. As a result, a multiplier of $C^{\ast}(\mathcal{G})$ is Fourier-transformed to a family of operators parametrized by the representation space $\widehat{\mathcal{G}}$, and the properties of this operator family reflect if the multiplier is an element of $C^{\ast}(\mathcal{G})$ itself or not. Explicitly, for an element $e\in\mathcal{M}(C^{\ast}(\mathcal{G}))$, if we can verify that the Fourier transform $\mathcal{F}(e)\in\mathcal{F}(\mathcal{M}(C^{\ast}(\mathcal{G})))$ actually lies in $\mathcal{F}(C^{\ast}(\mathcal{G}))$, then $e\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$. See the following diagram:
@@
For each $x\in M$, suppose that $Q$ is a multiplier of the $C^{\ast}(Gr(H)_{x})$ module, $Q\in\mathcal{M}(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$. Then, $Q$ is compact , i.e.,
@@
lies in the sub-algebra $\text{End}(E_{x})\otimes B^{\ast}(\widehat{Gr(H)}_{x})$. The algebra $B^{\ast}(S)$ will be defined in Definition 6.1.5.
An illustrative example of transversally H-elliptic operators is given by the operator $P_{\gamma}=\Delta_{n-k}+i\gamma T$ on the Heisenberg group $M=\mathbb{H}_{n}$ with transition action by $\mathbb{R}^{2k}$ on the first $2k$-coordinates. Here, $\Delta_{n-k}=-\sum_{i=k+1}^{n}({X_{i}}^{2}+Y_{i}^{2})$ means the sub-Laplacian operator on the sub-manifold $M^{\prime}\cong\mathbb{H}_{n-k}$ identified by coordinates $x_{k+1},y_{k+1},\dots,x_{n},y_{n},z$. The function $\gamma$ on $M$ is assigned such that $P_{\gamma}$ on $M^{\prime}$ is H-elliptic, see [baum2014k]. Utilizing Theorem 0.0.3, we can show that $P_{\gamma}$ satisfies the condition in Definition 0.0.1, so it is transversal H-elliptic on $M$.
@@
  • Chapter 3 articulates the general index theorem for H-elliptic operators, extending van Erp’s findings on contact manifolds [van2010atiyah1]. We reformulate this theorem from a $KK$-theoretic viewpoint.
@@
  • Chapter 5 is dedicated to the main index theorem for transversally H-elliptic operators. We will give a proof of the theorem using a similar approach as in [kasparov2022k] and [kasparov2024coarsepseudodifferentialcalculusindex].
  • Chapter 6 introduces the Fourier transform theory for $C^{\ast}$-algebra of nilpotent Lie groupoids [beltictua2017fourier] to verify the compactness condition in Definition 0.0.1.
  • In the Appendix, we exhibit some other possible conditions to define the transversally H-elliptic operators. We include prototypes of these definitions, from the definitions or certain crucial properties of transversally elliptic or H-elliptic operators.
@@
  • We denote by $G$ the locally compact Lie group acting on manifold $M$, by $\mathcal{G}$ a groupoid or a bundle of nilpotent Lie groups over the manifold $M$.
@@
  • If $E$ is a vector bundle over $M$, we will denote by $C^{\infty}(M;E)$ the space of smooth sections of $E$ and $C_{0}(M;E)$ the space of continuous sections of $E$ vanishing at infinity. We denote $C_{0}(E)$ to mean the space of continuous functions on $E$ (as a topological space) vanishing at infinity.
@@
  • We denote by $\mathcal{M}(A)$ the multiplier algebra for a $C^{\ast}$-algebra $A$.
@@
In [van2010atiyah1], van Erp analyzed the sub-elliptic operators on contact manifolds using the tangent groupoid approach and gave an index theorem for H-elliptic operators. The contact manifolds are special cases of filtered manifolds, and an analog of the tangent groupoid can be defined for filtered manifolds. We can extend the H-elliptic index theory to filtered manifolds.
In this chapter, we introduce some fundamental concepts related to filtered manifolds and the symbolic calculus of operators on filtered manifolds.
@@ 1.1 Filtered manifold and the osculating groupoid
Let $M$ be a smooth complete Riemannian manifold with a filtration of the tangent bundle
@@
Assume that $M$ is equipped with a proper isometric action by a compact group $G$.
@@
At each point $x$ on the filtered manifold $M$, there is a simply-connected nilpotent Lie group structure on $T_{x}M$, called the osculating group at $x$, denoted by $Gr(H)_{x}$.
@@
It can be proved that $[X,Y](x)$ mod $H^{i+j-1}_{x}$ depends only on $X(x)$ mod $H^{i-1}_{x}$ and $Y(x)$ mod $H^{j-1}_{x}$. Therefore, the Lie bracket structure over $gr(H)_{x}$ is well defined for each $x\in M$. The graded algebra $gr(H)_{x}$ at each point is called the osculating Lie algebra. The corresponding simply-connected nilpotent Lie group for each $x$ is the $Gr(H)_{x}$ we required. ∎
@@ 1.2 Kirillov’s orbit method
When dealing with the cosymbol on a filtered manifold, which will be defined in Section 1.4, it is necessary to replace the covectors $\xi\in T^{\ast}_{x}M$ in $\sigma(P,x,\xi)$ by unitary irreducible representations $\pi$ of $Gr(H)_{x}$.
As outlined in [kirillov2004lectures], Kirillov’s orbit method characterizes the unitary dual $\widehat{G}$ as a topological space and construct the unitary irreducible representations $\pi_{\Omega}$ associated with a coadjoint orbit $\Omega\in\mathfrak{g}^{\ast}$.
@@
Kirillov’s orbit method asserts that:
@@
This map from coadjoint orbits $\Omega\in\mathcal{O}(G)$ in $\mathfrak{g}^{\ast}$ to unitary irreducible representations of $\pi_{\Omega}\in\widehat{G}$ is a one-to-one correspondence.
@@ 1.3 Groupoid $C^{\ast}$-algebra
The space $Gr(H)$ defined in 1.1.3 is a groupoid. In place of the algebra $C_{0}(T^{\ast}M)$, we need to analyze the $C^{\ast}$-algebra $C^{\ast}(Gr(H))$ in the filtered manifold cases. In this section, we briefly introduce the groupoid $C^{\ast}$-algebra theory.
@@
      source and range maps $r,s:\mathcal{G}\rightrightarrows\mathcal{G}^{(0)}$;
@@
      a multiplication map $\mathcal{G}^{(2)}\to\mathcal{G}$, $(a,b)\mapsto ab$, where
@@
      an inverse map $\mathcal{G}\to\mathcal{G}$, $a\mapsto a^{-1}$, such that: